OpenStudy (anonymous):

plz! 6+x+3x-8+____=2x

OpenStudy (anonymous):

Combine like terms. -2+4x=2x -2=-2x x=2

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

wait where did you get -2x

OpenStudy (anonymous):

bringing the 4 to the right side of the equation makes it a -4x. -4x+2x =-2x

OpenStudy (anonymous):

ohh.. nevermind thank you